OpenStudy (anonymous):

solve using substitution: y=1/4x ; x=4(y+3)

OpenStudy (anonymous):

y=1/4x x=4y x=4(y+3) 4y=4(y+3) y=y+3 0=3 Those two equations will never equal each other

OpenStudy (anonymous):

no solution?

OpenStudy (anonymous):

These are the two lines created by the two equations You have y=(1/4)x and y=(1/4)x-3 Those two lines are parallel since they have the same slope. So no, no solution

OpenStudy (anonymous):

ahh, i see. thank you!